Consider the reaction 3CaCl2(aq) + 2Na3PO4(aq) --> Ca3(PO4)2(s) + 6 NaCl(aq) (molar mass of calcium phosphate is 310.18 g/mol)

What mass of Ca3(PO4)2 is produced from 94.2 mL of 0.250 M Na3PO4 solution?
1
Expert's answer
2020-04-06T11:47:21-0400

3CaCl2(aq) + 2Na3PO4(aq) -> Ca3(PO4)2(s) + 6NaCl(aq)

n(Na3PO4) = 0.25*0.0942 = 0.02355 mol

n(Ca3(PO4)2) = n(Na3PO4)/2 = 0.02355/2 = 0.011775 mol

m(Ca3(PO4)2) = 0.011775*310.18 = 3.6524 g